### v1

# The AI Memory Gap: Users Misremember What They Created With AI or Without

Abstract.
As large language models (LLMs) become embedded in interactive text generation, disclosure of AI as a source depends on people remembering which ideas or texts came from themselves and which were created with AI. We investigate how accurately people remember the source of content when using AI. In a pre-registered experiment, 184 participants generated and elaborated on ideas both unaided and with an LLM-based chatbot. One week later, they were asked to identify the source (noAI vs withAI) of these ideas and texts. Our findings reveal a significant gap in memory: After AI use, the odds of correct attribution dropped, with the steepest decline in mixed human-AI workflows, where either the idea or elaboration was created with AI. We validated our results using a computational model of source memory. Discussing broader implications, we highlight the importance of considering source confusion in the design and use of interactive text generation technologies.
Ideation, Writing, AI, LLM, Source Memory
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Natural language generation
Figure 1. Overview of our experiment and findings: In Phase 1 (left) participants were presented with problems and asked to come up with five ideas each (as 1-3 keywords per idea), with or without assistance by a chatbot (alternating across problems). They then wrote on these ideas (1 sentence per idea), again with or without AI. After a week, Phase 2 (centre) presented the problems and solutions again, asking participants if they worked on this (item memory), and if so, what the source was for both idea and elaboration (source memory). We also showed distractors (unseen problems/solutions). The key findings (right) show that AI had a negative impact, mixed workflows were particularly difficult, and people tended to be overconfident in their own performance.
[…]
Illustration showing the study flow overall: Left part shows Phase 1 with a user and chatbot working on ideas and elaborations for the example problem “How might we solve the problem of plastic waste in the ocean?” Text says that participants have to come up with five ideas in 1-3 keywords and write elaborations in one sentence each. An arrow with annotation (“One week later”) points to a second illustration of a user with question-marked though bubbles. Text says that in this Phase 2, participants were asked: Did you work on this? Source of idea? Source of elaboration? Further text says that they were also asked this for unseen items (so-called distractors). Following another arrow to the right is a box with the title “Key findings” and three bullets: Negative impact of AI on source memory overall; mixed workflows harder to remember than never/always using AI; and people tend to be overconfident about their own performance.
1. Introduction
Do you remember which ideas and sentences you authored yourself and which you created together with AI?
Co-creation with Large Language Models (LLMs) is rapidly evolving, as many people integrate AI into their daily workflows for a wide range of tasks (Mysore et al., 2025), with a growing body of research designing LLM-based systems for writing (Lee et al., 2024; Benharrak et al., 2024; Hoque et al., 2024), ideation (Shin et al., 2023; Suh et al., 2024; Shaer et al., 2024), developing concepts (Han et al., 2024; Wang et al., 2025), and solving complex problems (He et al., 2025a), which comes with metacognitive demands (Tankelevitch et al., 2024). However, despite the growing reliance on LLMs, little is known about people’s ability to accurately recall their own contributions in content co-created with AI.
As research shows, users generally perform at near-chance levels when trying to identify AI‑generated text (Clark et al., 2021; Jakesch et al., 2023b; Boutadjine et al., 2025), particularly non-regular users (Russell et al., 2025). Work with essay writing found that LLM users recalled less from their own texts and showed weaker EEG engagement than unaided or search users, though in a small sample (Kosmyna et al., 2025). However, there is no systematic study on how writers remember the sources of ideas and texts after working with AI.
At the same time, there is growing concern around authorship claims in AI-generated content (He et al., 2025b; Kyi et al., 2025), with evidence that users sometimes take credit for work largely or entirely produced by LLMs (Draxler et al., 2024).
This issue matters not only for questions of intellectual ownership, but also because LLMs can shape cognitive processes. For instance, prior research shows that LLMs can influence users’ opinions (Jakesch et al., 2023a), which suggests they might also affect how people remember their own ideas and contributions.
Decades of cognitive psychology indicate that when two sources produce highly similar linguistic output, especially within a single writing session, people rely on heuristic cues (e.g., ease of retrieval, writing fluency) rather than firm recollection to attribute authorship (Johnson et al., 1993). LLM co‑writing blurs those cues further because AI suggestions are (1) context‑specific, (2) immediately editable, and (3) increasingly integrated into the user’s document editing tools, making them potentially feel self-generated.
This new form of source confusion between self and AI could have important implications for perceived authorship, learning, accountability, and trust in both self and system. It may shape how users reflect on and take ownership of their work, particularly in educational, professional, or creative contexts. For instance, users could think less of their own expertise by discovering an AI-generated mistake and misremembering it as their own. Or they might treat a document as more reliable than warranted if they misremember AI-generated content as their own. Conversely, they may lose deserved “credit” in external evaluations by mistakenly disclosing an own idea as AI-generated.
However, we currently lack the fundamental understanding necessary to make informed design choices in this context. In particular, it remains unknown how accurately and confidently people can recall what they created themselves versus what was suggested by AI, and whether this depends on the AI’s role (e.g., idea generation, text composition, or both). This motivates our research questions:
  • RQ1: How does working with AI on a problem affect people’s ability and confidence in remembering (1) the solution, (2) the source of the idea, and (3) the source of its elaboration?
  • RQ2: How do mixed human-AI workflows (different sources for ideas and elaborations) compare to consistent workflows in terms of memory accuracy and confidence in source attribution?
  • RQ3: To what extent are people able to judge their own ability to remember whether an idea or elaboration originated from AI vs. themselves?
To address these questions, we conducted a within-subjects study with two phases (Figure 1). In Phase 1, participants (N=184) completed ideation and writing tasks either alone (noAI) or with an LLM-based chatbot (withAI). After a one-week delay (Phase 2), they were asked to recall the origin of each idea and elaboration sentence they had produced, specifying whether they had created it themselves or with AI support, and rating their confidence in each decision. To probe memory accuracy more rigorously, we also introduced content not previously seen by participants (“distractors”).
Our results reveal an “AI Memory Gap”, where any AI involvement impaired memory, specifically source memory, and most strongly in mixed human-AI workflows. The odds of correctly attributing the source of ideas were 95% lower when the idea originated from the AI but was elaborated by the human, and 86% lower when the idea was human but the elaboration was generated by the AI, compared to the all-human baseline. When participants consistently used AI in both steps, the interaction partly offset these losses. The same pattern holds true for elaboration source attribution, where using AI for the idea alone reduced the odds of correct attribution by 65%, and using AI for elaboration alone reduced them by 63%. When both idea and elaboration came from AI, the interaction again compensated for the losses, but still remained below the all-human baseline. Confidence mirrored these effects: participants were most confident in the all-human workflow and less confident with AI involvement. Yet, they systematically overestimated their performance, with self-rated accuracy exceeding actual performance by 12% for ideas, and 6% for elaborations.
These findings show that co-creating with LLMs can systematically impact source memory and attribution, and that the role of AI within the workflow matters in ways that open avenues for UI and interaction design.
Discussing broader implications, we highlight the importance of considering source confusion in the design of interactive text generation technologies.
2. Related Work
2.1. Ideation and Writing Assistants
We investigate source memory for both finding ideas and expressing them in text. Thus, here we relate our work to research on ideation with AI and writing with AI.
2.1.1. Ideation with AI
Recent related work addresses HCI aspects of integrating AI into ideation processes: For example, Shin et al. (2023) organised a CHI’23 workshop on collaborative ideation with AI and Han et al. (2024) investigated how pairs of students prompt ChatGPT and Midjourney to create stage design ideas. Similarly, Mahdavi Goloujeh et al. (2024) investigated users’ “prompting journeys” for text-to-image generation via interviews, extracting common structures and challenges. Their title asks “Is it AI or is it me?” as a reflection on creatives’ perception but not on remembering their creative actions.
Further work examined creatively stimulating effects of “AI errors”, such as mislabelled object recognitions (Van Der Burg et al., 2022). Liu et al. (2024) investigate such errors specifically for ideation, to promote human creativity through serendipity and association.
Other work reports controlled experiments on related (negative) effects: Guo et al. (2024) studied the impact of value alignment between user and chatbot in a brainstorming task (i.e. both “pro” or “con” on a topic – or in opposition). They found no difference in idea quality, but the generated ideas reflected the AI’s values and using AI reduced the feeling of ownership. Related, Wadinambiarachchi et al. (2024) found that novice designers using image generators had higher design fixation (i.e. keeping close to features of an initial example). They also produced fewer, less varied, and less original ideas with AI.
In contrast, avoiding fixation is a design goal in other work: CreativeConnect by Choi et al. (2024) helps graphic designers expand their space of ideas by extracting inspiring keywords from reference images on a moodboard. Similarly, AIdeation by Wang et al. (2025) supports idea exploration with image generation and keyword extraction in an iterative ideation loop. And Luminate by Suh et al. (2024) plots systematically varied alternatives on a 2D space of ideas.
While the above systems have different roles for what the user and AI add to the process and UI, the IdeationWeb system by Shen et al. (2025) used colour-coded nodes to mark otherwise similar representations of human and AI ideas, namely nodes in a mind map-like diagram. In contrast, we examine how people remember the source of ideas themselves.
In summary, there is high interest in applying AI in ideation tasks, with research on both constructive aspects (e.g. tool design) and assessing impact. Adding to this literature, we examine the new aspect of how people later remember the source of ideas when using AI in ideation. Specifically, our task involves both coming up with ideas and writing short descriptions, leading us to a review of writing with AI next.
2.1.2. Design and perception of writing with AI
The recent survey by Lee et al. (2024) analysed 115 “intelligent” writing tools from papers at HCI and NLP venues, charting a design space with five high-level dimensions. These design choices may impact how people remember their interactions and results. For example, their dimension of “Interaction” includes aspects such as UI layout (separated or integrated AI), interaction metaphor (e.g. tool or agent), and interface paradigm (e.g. chatbot or text editor). For our study, we decided to mimic the currently widespread, de-facto “standard UI” for LLMs: a chatbot in a sidebar.
Moreover, this survey (Lee et al., 2024) extracted aspects of the user’s long-term perception and relationship to the system (e.g. agency, ownership, integrity, trust, transparency). Memory was not discussed, reflecting the gap in the literature, although remembering what was created by AI or the user seems highly relevant for long-term perception. For instance, perceived ownership may be impacted by this, as well as integrity – for example, awareness of (accidental) plagiarism introduced by AI-generated content.
One notable exception is the HaLLMark system by Hoque et al. (2024), which is explicitly designed for writers to keep a history of prompts and AI contributions throughout their writing process. As with IdeationWeb (Shen et al., 2025) above, this tracks AI use in the UI, whereas we investigate how users remember the source of ideas and texts unaided.
2.2. Source Memory
Source memory refers to the processes that enable people to remember where, when, and how an item was acquired, in contrast to item memory, which concerns only whether the item was encountered at all (Johnson et al., 1993; Mitchell and Johnson, 2009). The dominant theoretical account is the Source Monitoring Framework (SMF), which posits that people attribute mental contents to a source by evaluating qualitative features of the memory trace – such as perceptual details, spatio-temporal context, associated affect, and records of the cognitive operations that generated the content (Mitchell and Johnson, 2009). When these features are insufficiently distinctive, individuals rely on heuristic cues such as familiarity or processing fluency, increasing the likelihood of misattribution.
Items that people generate themselves are typically recognised better than items they merely read or hear, a robust phenomenon known as the generation effect (Geghman and Multhaup, 2004). Paradoxically, generation can impair source memory under conditions of high similarity between self- and other-produced material. For instance, Landau and Marsh (1997) showed that when participants alternated between their own word puzzle solutions and copying sentences provided by an experimenter, they later misattributed large proportions of the material. Comparable effects – unknowingly claiming someone else’s ideas – have been reported for collaborative recall and brainstorming tasks (Brown and Murphy, 1989; Marsh et al., 1997), but not yet for “collaboration” with AI.
Digital environments introduce additional challenges for source monitoring. For instance, real-time collaboration and ubiquitous copy-paste affordances might blur perceptual and contextual cues that would otherwise separate one’s own contributions from external input. Code repository studies of end-user programmers, for example, document frequent misattribution of copied code snippets to self-authorship (Baltes and Diehl, 2019). In language generation interfaces, suggestions are inserted directly into the working document; users can accept, reject, or overwrite them at the keystroke level, intertwining human and machine text in ways that make the final product perceptually homogeneous (e.g. see (Buschek et al., 2021; Lee et al., 2022)). Under the SMF, such homogeneity should reduce cue distinctiveness and therefore impair source discrimination – a prediction supported by recent empirical work on AI-assisted writing (Draxler et al., 2024) and on detecting AI-generated texts (Boutadjine et al., 2025).
Confidence in source decisions is only weakly correlated with accuracy (e.g. see (Horry et al., 2014)). People consistently overestimate their ability to identify the origin of text, especially when they believe that stylistic idiosyncrasies are obvious cues such as in the case of AI writing (Fleckenstein et al., 2024). Metacognitive overconfidence is exacerbated by fluency – the ease with which information is processed – because fluent processing is misinterpreted as evidence of prior self-generation (Huang and Shanks, 2021). LLMs explicitly optimise for fluency, suggesting that AI collaboration may produce particularly compelling illusions of authorship.
Overall, despite a rich literature on source monitoring in laboratory tasks, little is known about source memory for content created in interactive human-AI settings.
3. Method
To investigate how co-writing with LLMs affects people’s source memory for ideas and elaboration texts, we conducted a two-phase, within-subjects online experiment. Figure 2 gives a visual overview.
Figure 2. Study design and procedure. The experiment consisted of two phases separated by one week. In Phase 1 (top), participants first generated five short ideas (1-3 keywords each) for eight problem statements, alternating between conditions with and without AI assistance (top left). Second, they elaborated on each idea with a one-sentence explanation, again alternating between AI-assisted and unassisted conditions (top right). Problem order and withAI/noAI order were counterbalanced across participants. In Phase 2 (bottom), participants saw 60 elaborations (40 self-generated and 20 distractors) together with the original problem statements. For each, they indicated whether they remembered working on this item, and if so, attributed the source of both the idea and the elaboration (self vs. AI), along with confidence ratings. Distractors included both elaborations from familiar problems (known-topic) and from unseen problems (unknown-topic). This procedure implemented a 2×2 within-subjects design (noAI vs. withAI during ideation and elaboration) to measure item memory and source attribution.
[…]
This figure presents a flow diagram of the study design. Phase 1 (top row): Participants are shown eight problems (out of a pool of ten). For each problem, they generate five short ideas (1–3 keywords each), sometimes with AI assistance and sometimes without. Icons of lightbulbs represent ideas, and robot icons indicate AI involvement. After ideation, participants elaborate on each idea by writing a one-sentence explanation, again alternating between AI-assisted and unassisted conditions. Problem order and AI/no-AI conditions are counterbalanced. Phase 2 (bottom row): After one week, participants complete a memory and source attribution task. They are shown elaborations linked to problems, including 40 items from Phase 1 and 20 distractors. Distractors come from both problems participants had seen before (known-topic) and problems they had not seen (unknown-topic). For each elaboration, participants indicate (1) whether they remember working on it, (2) whether the idea source was themselves or AI, and (3) whether the elaboration source was themselves or AI. They also provide confidence ratings. Legend: Blue icons represent participants’ own elaborations from Phase 1, gray icons represent distractors from previously seen problems, and pink icons represent distractors from unseen problems. Overall, the figure illustrates the two-phase procedure designed to test item memory and source attribution under conditions with and without AI support.
Participants (N=184) had to create ideas and text by themselves (noAI) or with the help of a chatbot (withAI) and were later asked to remember and indicate the sources.
The study was pre-registered. Details can be accessed at: https://aspredicted.org/g78p-w7db.pdf. Deviations from the pre-registration can be found in Table 9.
3.1. Experiment Design
Our experiment consists of two phases a week apart. In Phase 1, participants created solution ideas and elaboration texts, either with or without AI support. In Phase 2, they indicated the sources of idea and elaboration, given the problem and elaboration text. Participants who completed Phase 1 received an invitation after 6 days to complete Phase 2 within 48 hours. This delay is consistent with source memory studies in the literature (e.g., (Beaufort et al., 2013; Bornstein and Lecompte, 1995; Brown and Halliday, 1991)) and the 48-hour time window was used to reduce drop-out.
3.1.1. Independent Variables
The independent variables are the presence (withAI) or absence of AI chatbot assistance (noAI) during ideation and during elaboration, respectively, resulting in a 2×2 within-subjects factorial design. Each participant experienced all four experimental conditions:
  1. (1)
    no AI assistance during either ideation or elaboration,
  2. (2)
    AI-assisted ideation but manual elaboration,
  3. (3)
    manual ideation but AI-assisted elaboration, and
  4. (4)
    AI assistance during both ideation and elaboration.
Splitting this overall content creation task into the two subtasks (ideation and elaboration) allowed us to examine how the phase in which AI support is used affects source memory. The two subtasks also reflect common real-world use cases of LLMs: generating initial ideas (ideation) and expanding or refining ideas (elaboration).
In Phase 1, participants were asked to generate five ideas (1-3 keywords each) on how to solve diverse problems, alternating between conditions with and without AI support. We created ten different problem statements and each participant had to work on eight of those. The two remaining topics served as distractors in Phase 2. To control for order effects, we counterbalanced the ideation task across participants: if one participant began the ideation phase with AI support, the next began without it. We also counterbalanced the selection of the eight out of ten problems and their order of presentation.
Within Phase 1, after completing all ideation tasks, participants moved on to the elaboration step. Here, they were presented with their generated ideas in the same order in which they had created them. For each idea, they wrote a one-sentence elaboration, again alternating between the withAI and noAI conditions, and counterbalanced across participants.
3.1.2. Dependent Variables
We measured four main types of dependent variables:
  1. (1)
    item memory performance (i.e. remembering to have worked on a problem and solution at all or not)
  2. (2)
    source attribution accuracy for idea and elaboration
  3. (3)
    source attribution confidence for idea and elaboration
  4. (4)
    perceived task performance
In detail, Phase 2 displayed each participant’s own elaboration sentences along with the problem text (40 items). In the same manner, we displayed the distractors (texts not previously seen or generated by the participant, 20 items). For counterbalancing, we randomised the order of problems and solutions for each participant and used this in a “round-robin” procedure (one idea from all problems before returning to another idea from the first problem), thus maximizing the distance between encountering solutions for the same problem. This was motivated by reducing context effects and shortcuts in source attribution.
First, for each elaboration shown in Phase 2, participants indicated whether they remembered working on it (binary: “Yes” / “No”), allowing us to assess item memory performance. Second, they attributed the source of each item’s underlying idea and elaboration (binary: “On my own” / “With AI support”). Third, they rated their confidence in each attribution on a slider ranging from 0 (“Very unsure”) to 100 (“Very sure”).
Finally, at the end of Phase 2, participants estimated their own attribution accuracy for both idea and elaboration source (0–100%) and indicated whether they believed they had more often misattributed content to themselves or to the AI.
3.2. Participant Recruitment
Our sample comprises 184 generally healthy people. Out of those 89 self-identified as female and 95 as male. All participants were either based in the US (72 people) or the UK (112 people) and all of them listed English as their primary language, with all but 7 participants being native speakers. Their age ranged from 19 to 75 (mean=39.8, SD=12.7). We asked participants how often they use AI tools (e.g., ChatGPT, Claude, Copilot). Most of them (139) reported daily use or several times a week, some (35) once a week or several times a month, and few participants (10) rarely or never user AI. Most participants had a college degree: 82 held a bachelor’s, 28 a master’s, and 7 a doctorate. Eleven reported an associate degree. 37 had a high school diploma and 18 had some college education without a degree. One reported another type of education.
The sample was acquired through the online platform Prolific 111www.prolific.com. We excluded 16 participants from the initially recruited 200 people, as they either did not complete Phase 2 of the experiment within the required three-day window (6 people) or our logs showed that they did not complete the tasks correctly (10 people; e.g., entering “no idea” or always submitting the same text). Participants who completed both parts of the experiment received a total compensation of £23. The average completion time was 107 minutes, resulting in an effective hourly rate of approximately £12.90. The experiment was approved by our institution’s ethics committee.
3.3. Apparatus
Here, we describe the study materials for the tasks, followed by an overview of the experimental software.
3.3.1. Problem Statements
We prepared a pool of ten problem statements for the ideation task in Phase 1, spanning social, technological, and environmental issues. They were designed to spark creative problem-solving without triggering emotional responses or appearing overly opinionated (e.g., “List practical ways a mid-sized company could reduce employee susceptibility to phishing emails.”). Guided by the concept of desirable difficulty (Bjork and Bjork, 2020), we formulated the problems so that generating five ideas would be achievable for everyone within a reasonable amount of time, but challenging enough to encourage effortful cognitive processing. A list of all ten statements can be found in  Section A.1. For each participant, eight of those topics were used in Phase 1, and two served as distractors in Phase 2.
3.3.2. Distractors
To assess false memory and source misattribution, we included distractor items in Phase 2. These were texts that participants had not seen or generated during Phase 1. We implemented two types of distractors.
Known-topic distractors
For each problem statement a participant had encountered in Phase 1, we added one additional elaboration sentence from the same problem as a distractor. These known-topic distractors tested whether participants would falsely recognize or misattribute content that was topically familiar but otherwise novel. To reduce the risk that within-topic distractors might be inadvertently similar to participants’ own responses, which could lead to them “correctly” remembering distractors, we chose solution texts that are very distinctive and always included a proper name. For example, for the problem statement “Oceans are heavily polluted with plastic waste. What practical approaches could help reduce plastic pollution in the oceans?”, the within-topic distractor was: “Deploy AquaClean bio-magnetic kelp that’s genetically modified to attract and collect microplastics for easier harvesting.”
Unknown-topic distractors
We also constructed unknown-topic distractors: solution elaborations for problem statements participants had never worked on at all. For each of the ten possible problem statements in our pool, we created six elaboration texts: one was used as the known-topic distractor (described above), and five were rather obvious solutions that served as these unknown-topic distractors.
3.3.3. Experimental Software
We developed a custom web application using React  [^2] https://react.dev with Next.js  [^3] https://nextjs.org (Figure 3). It handled communication with a PostgresDB using Prisma [^4] https://www.prisma.io. We used this database to store participant responses as well as some interaction logs. We used the AI-SDK [^5] https://ai-sdk.dev to enable communication between users and the LLM endpoint. We used the OpenAI endpoint “gpt-4.1-mini-2025-04-14” with the system prompt “You are a helpful assistant.” and provided no further context.
For Phase 1, the app showed a welcome screen with study information and consent form, followed by the tasks (Figure 3). In Phase 2, it then displayed the corresponding memory and attribution tasks (Figure 4).
Figure 3. Creating ideas and elaborations in Phase 1: (A) In the ideation phase participants entered five ideas for each problem into the respective input fields. (B) After creating ideas for every problem, they elaborated on those ideas in one sentence each. AI-support alternated for each problem in the ideation task and each idea in the elaboration task. Therefore all participants created half of the ideas and elaborations with AI support (C) and the other half without (D). The progress bar (E) indicates how far along participants are in the study.
[…]
The figure shows a multi-panel screenshot of a web-based study interface. Panel A depicts the ideation task with input boxes for five ideas, a problem statement, and a chat panel on the right where an AI assistant provides keyword suggestions. Panel B shows the elaboration task where participants write one-sentence elaborations for each idea. Panels C and D illustrate the AI chat area, either active with AI-generated keywords (C) or locked without AI support (D). Panel E displays a progress bar at the bottom indicating 10% completion of the study. Arrows and labels connect the panels to clarify workflow and indicate alternating AI support across tasks.
Figure 4. Memory and source attribution in Phase 2: We displayed the task description, the original problem statement, and the solution text created in Phase 1. Participants first had to answer whether they remember working on a solution, if so they had to indicate whether they came up with the underlying idea and the solution text on their own or with AI support. They also had to provide confidence ratings (0-100) for each.
[…]
The figure presents a web-based recognition task interface. At the top, a blue box explains the task instructions, followed by a green box showing the problem statement and a purple box displaying the solution text produced earlier. Below, participants respond to questions: a Yes/No toggle asks if they remember working on the solution, and two sets of radio buttons and sliders record whether the underlying idea and the written solution were generated independently or with AI support, along with confidence ratings from 0 to 100. The layout uses distinct colored sections and interactive elements for clarity.
3.4. Procedure
All participants completed the study online via Prolific, using a web browser on a desktop or laptop computer. Mobile devices were not permitted. Participants first read an online information sheet and consent form, and if they agreed continued to Phase 1. On average, participants required 74.7 minutes (SD 30.9 minutes) for Phase 1 and 32.8 minutes (SD 16.6 minutes) for Phase 2. After completing Phase 1, they started Phase 2 after 6 days and 5 hours (SD 8 hours 46 minutes) on average.
3.4.1. Phase 1 – Creating Ideas and Elaborations
The task description displayed the problem text and instructed participants to create five ideas each consisting of 1-3 keywords. These ideas were entered into five input fields displayed on the same page (Figure 3 A).
In the withAI condition, a chatbot interface was available on the right side of the screen (Figure 3 C). Participants could copy text from the chatbot and paste it into the input fields using the usual keyboard shortcuts or buttons located next to each field. In the noAI condition, the chatbot was locked (Figure 3 D) and pasting into the input fields was disabled.
After entering keywords into all five fields, the “Next” button became active, allowing participants to move on to the next ideation task. Each participant completed a total of eight such tasks, alternating between conditions with and without AI support, resulting in a total of 40 ideas.
Following this, participants were asked to elaborate on each of their previously submitted ideas with a one-sentence elaboration. Each such elaboration task displayed the task instruction, the original problem statement, and one of the participant’s earlier ideas alongside a text input field (Figure 3 B).
Each elaboration page featured one idea at a time, in the order they were submitted. Participants completed these elaborations with and without AI assistance in alternating order.
For both ideation and elaboration, the overall progress was shown at bottom of the application (Figure 3 E). After completing elaborations for all 40 ideas, the study concluded with a thank-you screen. Participants were reminded that they would receive an invitation to Phase 2 of the study six days later and that this follow-up phase would then need to be completed within 48 hours.
3.4.2. Phase 2 - Memory and Source Attribution
After the six-day delay, participants who completed Phase 1 received an invitation to take part in Phase 2. They were presented with the study information and were again required to provide informed consent before proceeding.
We first asked participants to guess the purpose of the study in a free-text entry field. The main part of Phase 2 then consisted of the memory and source attribution tasks (Figure 4). Besides the overall task description, each task page displayed the original problem statement and a solution (elaboration) text created by the participant in Phase 1. The underlying ideas were not shown.
Participants first indicated whether they remembered working on the displayed solution at all. If they did, they were asked: (1) Whether they came up with the underlying idea on their own or with AI support, (2) How confident they were in that attribution, (3) Whether they wrote the solution text on their own or with AI support, and (4) How confident they were in that second attribution. If they did not remember working on the displayed solution at all, they skipped the attribution questions and proceeded to the next memory task.
This process was repeated for all 40 elaborated ideas from Phase 1, as well as for 8 distractor ideas related to the same problems participants had worked on (known-topic distractors), and 12 distractor ideas from two entirely new problems (unknown-topic distractors), for a total of 60.
At the end, participants filled out a final questionnaire. They were asked to: (1) Estimate what percentage of ideas they believed they had correctly attributed, (2) Indicate whether they more often misattributed ideas to themselves or to the AI, (3) Estimate what percentage of solution texts they believed they had correctly attributed, and (4) Indicate whether they more often misattributed texts to themselves or to the AI.
Additionally, participants were asked to explain any strategies they used to determine the source of each idea or solution text.
Finally, participants completed a demographic questionnaire, including questions about gender, English proficiency, education level, and frequency of AI tool use. An overview of all questions and answer options can be found in Section A.2.
Upon finishing both parts of the study, participants were shown a final thank-you screen.
3.5. Data sharing
We make our experimental software, study material, analysis code, and the collected data accessible through our project repository: https://osf.io/anonymous/. Two researchers went through all the data collected via free text input fields to ensure that all potentially privacy-sensitive information were removed before publication.
4. Results
We first examine people’s ability and confidence in remembering content sources when working with AI. Next, we assess participants’ perceived performance. To then disentangle underlying cognitive processes, we explore a Multinomial Processing Tree (MPT) model, allowing us to estimate separate parameters for memory, guessing, and bias in source attribution. Finally, we report on participants’ qualitative feedback to further contextualise our findings.
4.1. Source Attribution Accuracy
We computed binomial generalized linear mixed models (GLLMs) in R (R Core Team, 2020), using the packages lme4 (Bates et al., 2015) and lmerTest (Kuznetsova et al., 2017), with participants (ID) as random intercepts to test how the (ground-truth (GT)) sources of the idea and elaboration affected three types of memory: whether participants remembered (a) working on the solution at all, (b) the source of the idea, and (c) the source of the elaboration.
Figure 5 shows the model-based interaction plots of predicted source attribution accuracy for all three memory types across conditions (noAI vs. withAI).
In this section, the odds ratios (ORs) obtained from the GLMMs represent the effect sizes, quantifying how strongly the source factors influenced each type of memory.
Figure 5. Model-based interaction plots of predicted source attribution accuracy for three types of memory: whether participants remembered working on the solution, remembered the source of the idea, and remembered the source of the elaboration. Probabilities are estimated from binomial GLMMs with participants as random intercepts. Error bars represent 95% confidence intervals. The plots show that item memory was high across conditions, while idea and elaboration source memory exhibited strong interactions: Ideas created by participants without AI were better remembered when the elaboration was also created without AI. In contrast, ideas generated with AI were better remembered when AI was also used to elaborate on them.
[…]
This figure shows the model-based interaction plots of predicted recognition accuracy across human- and AI-sourced ideas and elaborations. The left panel shows that participants almost always remembered working on the solution, regardless of source. In contrast, the middle and right panels reveal clear interaction effects: participants more accurately remembered the source of the idea and elaboration when the elaboration’s source matched the idea’s source (noAI – noAI or withAI – withAI), but accuracy dropped when the sources were mixed.
4.1.1. Remembering the Item
[…]
Both main effects were negative and significant, and so was the interaction: $\beta_{\text{idea({withAI}{})}}=-0.50$ ($p<.001$), $\beta_{\text{elab({withAI}{})}}=-1.19$ ($p<.001$), $\beta_{\text{idea({withAI}{})}\times\text{elab({withAI}{})}}=-0.41$ ($p=.024$). In odds-ratio terms, using AI for the idea reduced the odds of correctly remembering that one had worked on the solution (OR =0.61=0.61), as did using AI for elaboration (OR =0.30=0.30); the combined effect of AI at both stages further reduced the odds (OR =0.67=0.67).
Table 1 shows the model-based predicted probabilities (from emmeans, response scale):
Idea
source (GT)
Elaboration source (GT) Pr(correct) [95% CI]
noAI noAI 97.6% [97.6, 98.8]
withAI noAI 97.3% [96.1, 98.1]
noAI withAI 94.7% [92.7, 96.2]
withAI withAI 87.9% [84.0, 90.9]
Table 1. Model-based predicted probabilities for correctly remembering having worked on a solution at all.
[…]
The table lists model-based predicted probabilities of participants correctly remembering that they worked on a solution. Four rows represent combinations of idea source and elaboration source, either without AI (noAI) or with AI (withAI).
Pairwise contrasts (Holm-adjusted) showed that the no-AI workflow (noAI/noAI) yielded higher accuracy than any condition involving AI: vs. withAI/noAI, OR =2.02=2.02; vs. noAI/withAI, OR =3.29=3.29; vs. withAI/withAI, OR =8.15=8.15, all $p<.001$. Among the AI-involved conditions, withAI/noAI exceeded noAI/withAI (OR =1.99=1.99) and withAI/withAI (OR =4.01=4.01); noAI/withAI also exceeded withAI/withAI (OR =2.45=2.45); again all $p<.001$. These results indicate that both AI idea generation and AI elaboration independently impair memory, and their joint use compounds this effect, with the strongest decrement observed when AI contributed at the elaboration stage.
4.1.2. Remembering the Source of Ideas
[…]
Both main effects were negative and significant, qualified by a strong positive interaction: $\beta_{\text{idea({withAI}{})}}=-3.05$, $\beta_{\text{elab({withAI}{})}}=-1.97$, $\beta_{\text{idea({withAI}{})}\times\text{elab({withAI}{})}}=3.81$; all $p<.001$. That is, using AI for the idea alone sharply reduced correct attribution (OR =0.05=0.05), as did using AI only for elaboration (OR =0.14=0.14); when both idea and elaboration came from AI, the interaction increased the odds ($\times 45.3$), partially offsetting the main-effect penalties.
Model-based predicted probabilities (Table 2) make the pattern clear:
Idea
source (GT)
Elaboration source (GT) Pr(correct) [95% CI]
noAI noAI 92.4% [91.5, 93.2]
withAI noAI 37.7% [33.6, 41.9]
noAI withAI 64.0% [60.2, 67.6]
withAI withAI 79.3% [75.9, 82.4]
Table 2. Model-based predicted probabilities for correctly remembering the idea source.
[…]
The table presents predicted probabilities of participants correctly recalling whether the idea was generated with or without AI support. Four rows represent combinations of idea source and elaboration source, either without AI (noAI) or with AI (withAI).
Participants were most accurate when no AI was involved at all (noAI/noAI), and least accurate when the idea came from AI but participants elaborated themselves (withAI/noAI). Notably, keeping the workflow consistent (withAI/withAI) substantially improved accuracy relative to the two mixed workflows (79.3% vs. 37.7% and 64.0%), indicating that mismatched sources between ideation and elaboration particularly impair memory for where the idea originated.
4.1.3. Remembering the Source of Elaborations
[…]
Both main effects were negative and significant, qualified by a strong positive interaction: $\beta_{\text{idea({withAI}{})}}=-1.05$, $\beta_{\text{elab({withAI}{})}}=-0.99$, $\beta_{\text{idea({withAI}{})}\times\text{elab({withAI}{})}}=1.77$; all $p<.001$. As such, using AI for the idea alone reduced correct elaboration-source attribution (OR =0.35=0.35), as did using AI only for elaboration (OR =0.37=0.37); when AI was used for both idea and elaboration, the interaction increased the odds ($\times 5.90$), partially compensating for the main effects.
Model-based predicted probabilities (Table 3) were:
Idea
source (GT)
Elaboration source (GT) Pr(correct) [95% CI]
noAI noAI 91.5% [90.1, 92.7]
withAI noAI 79.0% [76.3, 81.5]
noAI withAI 80.0% [77.3, 82.5]
withAI withAI 89.2% [87.2, 91.0]
Table 3. Model-based predicted probabilities for correctly remembering the elaboration source.
[…]
The table presents predicted probabilities of participants correctly recalling whether the elaboration was generated with or without AI support. Four rows represent combinations of idea source and elaboration source, either without AI (noAI) or with AI (withAI).
Participants were most accurate when no AI was involved at all (noAI/noAI) or when both idea and elaboration came from AI (withAI/withAI), with accuracy dropping in mixed workflows. Accuracy was lowest when AI provided the idea but participants elaborated themselves (withAI/noAI) or when participants provided the idea but AI elaborated (noAI/withAI).
4.1.4. Summary
Across all three memory measures, AI involvement reduced accuracy, though the magnitude and nature of this effect differed by memory type:
For remembering having worked on the solution at all (item memory), performance was generally high yet AI use lowered the odds of correct recognition by about 40% for ideation and by about 70% for elaboration.
For remembering the idea source, accuracy was substantially lower. Participants were most accurate when no AI was involved. Using AI consistently was less accurate (withAI/withAI) but higher than mixed workflows: Using AI for only the idea or only the elaboration reduced the odds, with withAI/noAI being the least accurate.
For remembering the elaboration source, the ranking of conditions was identical to idea-source memory, but overall accuracy was higher.
4.2. Source Attribution Confidence
We used LMMs with participant ID as a random intercept to examine how the ground-truth (GT) sources of ideas and elaborations shaped participants’ confidence in their attributions. Analyses focused separately on (a) confidence in remembering the idea source and (b) confidence in remembering the elaboration source.
Afterwards, we investigate the effect of correctness on confidence, by looking at the pairwise contrasts of correct vs. incorrect attributions, to examine participants’ metacognitive sensitivity.
Lastly, we use Pearson point-biserial correlation to examine how participants’ confidence correlates with their source attribution accuracy.
As confidence was measured on a continuous 0–100 scale, we report model-based estimated marginal means (EMMs) and regression coefficients ($\beta$) rather than odds ratios. These values serve as effect sizes in this section, quantifying how strongly the source factors influenced participants’ confidence in their attributions.
As an overview, Figure 6 shows interaction plots of model-based mean confidence in idea and elaboration attribution.
Figure 6. Confidence by idea and elaboration ground truth sources. Points show estimated marginal means; error bars show adjusted 95% CIs.
[…]
The figure is a two-panel line plot displaying participants’ confidence (0–100) in remembering idea and elaboration sources. The left panel shows confidence for the ground-truth idea source, and the right panel for the ground-truth elaboration source. Each panel plots two lines: a green line with circular markers for elaborations created without AI (noAI) and a red line with triangular markers for elaborations created with AI (withAI). Confidence is higher when both idea and elaboration are noAI, and lower when the idea is withAI. Error bars represent adjusted 95% confidence intervals.
4.2.1. Confidence in Remembering the Idea Source
For confidence in remembering the idea source, the grand mean across conditions was 81.10 (95% CI [79.29, 82.91]). Confidence was highest when both idea and elaboration originated without AI support (EMM = 85.2). Confidence dropped when AI was involved in either stage (EMMs = 80.0 and 80.3) and was lowest when AI was used in both stages (78.9). A significant Idea $\times$ Elaboration interaction ($\beta=0.95$, $p<.001$) indicated that the reduction with AI involvement was not purely additive. Figure 6 (left) shows the interaction.
The model-based estimated marginal means (Table 4, from emmeans, 0–100 scale) were:
Idea
source (GT)
Elaboration
source (GT)
Confidence [95% CI]
noAI noAI 85.2 [82.8, 87.6]
withAI noAI 80.0 [77.5, 82.5]
noAI withAI 80.3 [77.8, 82.9]
withAI withAI 78.9 [76.4, 81.3]
Table 4. Model-based estimated marginal means for confidence in remembering the idea source.
[…]
The table reports model-based estimated marginal means for participants’ confidence in remembering the idea source, with values on a 0–100 scale. Four rows represent combinations of idea source and elaboration source, either without AI (noAI) or with AI (withAI).
Pairwise contrasts (Holm-adjusted) showed that the no-AI workflow (noAI/noAI) yielded significantly higher confidence than any AI-involved condition ($p<.001$). Among those, differences were generally small, with only a modest advantage of noAI/withAI over withAI/withAI ($p=.024$).
4.2.2. Confidence in Remembering the Elaboration Source
For elaboration source confidence, the grand mean across conditions was 84.0 (95% CI [82.2, 85.9]). Confidence was highest when both idea and elaboration were created without AI (EMM = 86.3). It dropped when AI was involved at either stage: ideas ($\beta=-0.82$, $p<.001$) and elaborations ($\beta=-0.63$, $p<.001$). A significant interaction ($\beta=0.81$, $p<.001$) indicated that when both stages were AI, confidence was higher than expected from adding the two main effects. Figure 6 (right) shows the interaction.
The model-based estimated marginal means (Table 5) were:
Idea
source (GT)
Elaboration
source (GT)
Confidence [95% CI]
noAI noAI 86.3 [83.8, 88.7]
withAI noAI 83.0 [80.5, 85.5]
noAI withAI 83.4 [80.9, 85.9]
withAI withAI 83.4 [80.9, 85.8]
Table 5. Model-based estimated marginal means for confidence in remembering the elaboration source.
[…]
The table reports model-based estimated marginal means for participants’ confidence in remembering the elaboration source, with values on a 0–100 scale. Four rows represent combinations of idea source and elaboration source, either without AI (noAI) or with AI (withAI).
Pairwise contrasts (Holm-adjusted) confirmed that the all-human workflow (noAI/noAI) elicited significantly greater confidence than any AI-involved workflow ($p<.001$). By contrast, we found not significant differences between the AI-involved workflows.
4.2.3. Metacognitive Sensitivity
Metacognitive sensitivity describes the ability to accurately evaluate one’s own decisions, thoughts, or perceptions by distinguishing between correct and incorrect responses. We operationalized this as the difference in confidence between (ground truth) correct and incorrect attributions (Correct - Incorrect), estimated with emmeans from the integrated models that included ground truth and correctness factors. Positive values indicate higher confidence when the attribution was correct.
Idea source confidence
Metacognitive sensitivity varied strongly by GT combination, as can be seen in Table 6.
Idea
source (GT)
Elaboration
source (GT)
Idea $\Delta$ [95% CI] $p$
noAI noAI 10.57 [7.79, 13.34] ¡.0001
withAI withAI 5.04 [2.60, 7.48] .0001
noAI withAI 1.65 [-1.39, 4.70] .288
withAI noAI 0.83 [-1.43, 3.09] .472
[…]
The table reports model-based estimates of metacognitive sensitivity (delta = Correct − Incorrect) in confidence-scale points. Four rows represent combinations of idea source and elaboration source, either without AI (noAI) or with AI (withAI).
Pairwise comparisons (Holm-adjusted) showed that noAI/noAI exhibited a larger sensitivity than withAI/noAI (difference in $\Delta=9.74$, $p<.0001$), noAI/withAI ($8.92$, $p=.0001$), and withAI/withAI ($5.53$, $p=.0136$); withAI/withAI exceeded withAI/noAI ($4.21$, $p=.0366$); other differences were not significant. Thus, participants’ confidence tracked correctness most strongly without AI, moderately when using AI for both idea and elaboration, and weakly in mixed workflows. Figure 7 shows the confidence for Correct vs. Incorrect attributions within each ground truth combination.
Figure 7. Metacognitive sensitivity plots for idea: confidence for Correct vs. Incorrect attributions within each ground truth combination. Points are estimated marginal means with the same (Holm-adjusted) 95% CIs as used elsewhere; green = Correct, red = Incorrect. idea-source confidence shows a large correctness gap in noAI/noAI, a moderate gap in withAI/withAI, and minimal gaps in mixed workflows.
[…]
The figure displays a line graph of metacognitive sensitivity for idea confidence, comparing correct versus incorrect source attributions across four ground-truth idea–elaboration combinations. The x-axis lists the combinations (noAI × noAI, withAI × noAI, noAI × withAI, withAI × withAI), and the y-axis shows confidence in idea-source attribution (0–100). Green circles represent correct attributions and red crosses represent incorrect ones, with vertical error bars showing confidence intervals. The largest confidence gap between correct and incorrect occurs for noAI × noAI, a moderate gap appears for withAI × withAI, and minimal gaps are visible in the mixed-source conditions.
Elaboration source confidence
Here, metacognitive sensitivity was positive in all cases, as can be seen in Table 7.
Idea
source (GT)
Elaboration
source (GT)
Elaboration $\Delta$ [95% CI] $p$
noAI noAI 5.54 [2.99, 8.09] ¡.0001
withAI withAI 10.95 [8.70, 13.20] ¡.0001
noAI withAI 6.47 [3.66, 9.28] ¡.0001
withAI noAI 7.79 [5.71, 9.88] ¡.0001
[…]
The table reports model-based estimates of metacognitive sensitivity (delta = Correct − Incorrect) in confidence-scale points. Four rows represent combinations of idea source and elaboration source, either without AI (noAI) or with AI (withAI).
Contrasts (Holm) indicated a larger sensitivity in withAI/withAI than noAI/noAI (difference in $\Delta=5.41$, $p=.0112$); the remaining differences among AI-involved workflows were not significant (all $p>.17$). Thus, participants’ confidence in elaboration attribution tracked correctness in all workflows, with the highest sensitivity when AI was used for both idea and elaboration. Figure 8 shows the confidence for Correct vs. Incorrect attributions within each ground truth combination.
Figure 8. Metacognitive sensitivity plots for elaboration: confidence for Correct vs. Incorrect attributions within each ground truth combination. Points are estimated marginal means with the same (Holm-adjusted) 95% CIs as used elsewhere; green = Correct, red = Incorrect. elaboration-source confidence shows positive correctness gaps in all cells, largest in withAI/withAI.
[…]
The figure is a line plot of metacognitive sensitivity for elaboration confidence, comparing correct versus incorrect source attributions across four ground-truth idea–elaboration combinations. The x-axis shows combinations (noAI × noAI, withAI × noAI, noAI × withAI, withAI × withAI), and the y-axis represents confidence in elaboration-source attribution from 0 to 100. Green circles with error bars indicate correct attributions and red crosses with error bars show incorrect attributions, with confidence intervals. All conditions display a positive gap where correct attributions receive higher confidence, with the largest difference in the withAI × withAI condition.
4.2.4. Confidence-Accuracy Correlation
We quantified metacognitive calibration by correlating trial-level confidence with attribution accuracy (Correct =1=1, Incorrect =0=0) using Pearson point-biserial correlations. We report overall effects and effects within each ground truth combination of idea and elaboration sources (all $p$-values after Holm correction).
Overall calibration
Confidence reliably tracked correctness for both outcomes, with stronger coupling for elaborations than ideas ($r_{\text{idea}}=.115$; $r_{\text{elab}}=.218$; both $p<.001$). Thus, participants’ confidence was a small but dependable indicator of being correct, especially for elaboration-source judgments.
By ground-truth combination
For idea judgments, calibration depended on the workflow: Confidence tracked accuracy in consistent workflows (noAI/noAI $r=.228$; withAI/withAI $r=.234$; both $p<.001)$, but not in mixed workflows: noAI/withAI was essentially null (r=.001r=-.001, $p=.976$), and withAI/noAI showed a small negative association (r=.094r=-.094, $p=.0011$).
For elaboration judgments, confidence-accuracy coupling was positive and significant in all cases. It was strongest for consistent workflows, particularly withAI/withAI ($r=.277$), and more modest in mixed workflows (noAI/withAI: $r=.169$ ; withAI/noAI: $r=.192$).
Table 8 gives an overview.
Outcome (GT) Idea $\times$
Elaboration
$\mathbf{r}$ [95% CI] $p$
Idea noAI/noAI .228 [.186, .268] <.001<.001
Idea withAI/withAI .234 [.189, .278] <.001<.001
Idea noAI/withAI $\approx 0$ [.056-.056, .054] .976
Idea withAI/noAI .094-.094 [.147-.147, .041-.041] .0011
Elaboration noAI/noAI .215 [.173, .256] <.001<.001
Elaboration withAI/withAI .277 [.232, .320] <.001<.001
Elaboration noAI/withAI .169 [.115, .222] <.001<.001
Elaboration withAI/noAI .192 [.140, .243] <.001<.001
Table 8. Point-biserial correlations between confidence and accuracy within each GT combination. Holm-adjusted $p$-values control the familywise error rate within each outcome.
[…]
The table reports point-biserial correlations between confidence and accuracy within each ground-truth idea–elaboration combination. Four rows per outcome (Idea and Elaboration) represent combinations of idea source and elaboration source, either without AI (noAI) or with AI (withAI), for a total of eight rows.
4.2.5. Summary
AI involvement reduced confidence in source attributions, with the highest confidence in all-human workflows.
For idea-source confidence, consistency mattered: human-only workflows were highest, fully AI-supported workflows lowest, and mixed workflows in between. For elaboration-source confidence, all AI-involved workflows showed similarly reduced confidence.
Correctness boosted confidence (metacognitive sensitivity), but patterns differed: strong for ideas created without AI, moderate with full AI-support, and negligible in mixed workflows; for elaborations, sensitivity was positive in all workflows, largest with full AI.
Overall, confidence was a reliable but modest predictor of accuracy, with stronger calibration for elaborations than ideas.
4.3. Objective vs. Perceived Performance
For this analysis, we first computed per-participant accuracy rates for both idea source and elaboration source attribution as follows: For items seen in Phase 1, a response was correct if the item was recognised as seen and the reported source matched the ground truth. For distractor items, a response was correct if participants correctly detected that they had not worked on this in Phase 1.
We then compared these objective performance rates to participants’ self-assessments, collected after the recognition tasks. We used Spearman rank correlations to assess the correlation between actual and perceived performance. Because the self-assessments were expressed as percentages (0-100%), we rescaled them to the 0-60 range used for the objective measure (i.e., number of correct attributions out of the max of 60) for comparability.
Correlations between objective and perceived performance
For ideas, the association between source attribution accuracy and self-assessment was weak and not significant, $\rho=.13$ ($p=.077$). By contrast, it was significant for elaborations, $\rho=.25$ ($p<.001$).
Calibration ratios
Calibration ratios (self-rated / objective performance) exceeded 1 for both tasks, indicating general overconfidence: idea, $M=1.12$, $t(183)=3.81$ ($p<.001$); elaboration, $M=1.06$, $t(183)=2.03$ ($p=.043$). Ratios were significantly higher for idea than elaboration, $t(183)=3.22$ ($p=.0015$).
Figure 9 plots objective accuracy against participants’ self-reports for each measure.
Figure 9. Relationship between objective accuracy and self-assessed performance for idea- and elaboration-source attribution. The dashed line represents perfect calibration.
[…]
The figure is a scatter plot comparing objective performance (0–60) on the x-axis with self-assessed performance (0–60) on the y-axis for idea- and elaboration-source attribution. Yellow circles represent idea data and blue squares represent elaboration data. Solid yellow and blue regression lines show the relationship for idea and elaboration, respectively. A black dashed diagonal line indicates perfect calibration where self-assessment equals objective performance. Both regression lines have positive slopes, indicating higher self-assessments with higher objective scores, though many points lie above the dashed line, suggesting overestimation.
Together, these findings suggest that participants’ metacognitive monitoring of source attribution performance was more accurate for elaborations than for ideas. Elaboration self-assessments almost tracked actual performance and showed minimal bias, whereas idea self-assessments were weakly related to performance and systematically overestimated accuracy. Across both tasks, participants showed a tendency toward overestimating their performance, more strongly for ideas than elaborations.
4.4. Multinomial Processing Tree (MPT) Model
Recognition experiments with categorical responses are better analysed with MPT models (Bröder and Meiser, 2007). For a tutorial, see Schmidt et al. (2023). An MPT specifies a small set of latent cognitive states (e.g. “detect old”, “remember source”, “guess”) and the probabilistic paths that link those states to the observed response frequencies. MPTs therefore serve as computational models of recognition memory. Compared with raw proportion tests or LMMs, MPTs disentangle cognitive memory processes from pure response and yield parameter estimates that are directly interpretable in psychological terms.
Because joint trees of considering idea and elaboration recognition together failed to reproduce the strong within-item covariance, we modelled idea and elaboration responses with two independent 2-high-threshold (2-HT) MPTs (Snodgrass and Corwin, 1988). The component-wise 2-high-threshold model assigns a single probability to each latent processing step that links the unobserved memory state to the observed categorical response. All parameters are constrained to the unit interval and are estimated hierarchically (i.e., a group mean and an across-person standard deviation). For every component, the four free parameters have the following psychological interpretation, as shown in Figure 10.
All parameters were estimated hierarchically with TreeBUGS (Heck et al., 2018) using 40 000 MCMC iterations (after 5 000 adaptation), three parallel chains, and a thinning interval of 10. TreeBUGS fits MPT in a Bayesian hierarchical framework, allowing the researcher to account for repeated measures and multiple subjects, sampling the joint posterior of individual and group parameters with JAGS and returning full uncertainty intervals plus posterior-predictive $p$-values for mean and covariance structure. We used standard priors to facilitate convergence. Gelman–Rubin diagnostics were excellent ($\hat{R}\leq 1.01$). Posterior-predictive checks showed small mean discrepancies (T1 obs–pred <5<5), but covariance misfit (T2, $p=0$) indicates residual trial-level dependence that separate trees cannot capture. Parameter estimates are nevertheless robustly addressed for our preregistered H1.
Figure 10. Overview of the MPT model: After an item appears at test, $d$ is the probability that the participant recognises it as old. A higher $d$ therefore reflects greater mnemonic sensitivity, independent of any knowledge about the item’s source. When recognition of an old item fails (1d1-d), the model routes the trial directly to the Miss category. Note that the earlier GLMM outcome “remembered working on it” is not identical to the MPT detection parameter $d$: whereas $d$ represents a latent probability of correctly detecting studied items, the GLMM measure pools across hits, misses, and distractor rejections, so the two indices are not directly comparable. When the source is forgotten, the model assumes the participant guesses. Parameter $\beta$ is the probability that this guess is “withAI”; $1-\beta$ is therefore the probability of guessing “noAI”. Values of $\beta$ above $0.5$ indicate a systematic bias toward attributing the item to the AI system, whereas $\beta=0.5$ corresponds to an unbiased coin flip. On distractor trials there is, by definition, no true episodic trace. Parameter $f$ is the probability of producing an old response nevertheless. The ensuing attribution is, again, governed by $\beta$, yielding the False Alarm (withAI) or False Alarm (noAI) categories. The complement 1f1-f leads to the category Correct Rejection. Hence, $f$ quantifies the overall response criterion, with higher values signalling a more liberal “old” bias.
[…]
The figure presents a tree diagram explaining the Multinomial Processing Tree (MPT) model for source recognition. It begins with a test item that is either actually “old” or “new.” For old items, recognition occurs with probability d; if not recognized (1–d), the path leads to a Miss. Recognized items proceed to a Source decision with probability s. Correct source recall results in a Hit Correct outcome depending on whether the true source is “withAI” or “noAI.” If the source is not recalled (1–s), the model includes a guess step with probability beta for guessing “withAI” and 1–beta for guessing “noAI,” producing either Hit Correct or Hit Incorrect outcomes. For new items, a response of “old” occurs with probability f, leading to a guess step and possible False Alarm (withAI or noAI); a “new” response (1–f) is a Correct Rejection. Colored labels distinguish MPT model parameters (blue), incorrect outcomes (red), and correct outcomes (green). Text below the diagram describes each parameter and its interpretation in the model.
4.4.1. Idea component
For ideas, the model estimated a hit rate of $d_{\text{idea}}=.86$ (95% HDI $[.84,.88]$), indicating that the vast majority of studied ideas were correctly recognised as old. Conditional on a hit, the true source was recalled on $s_{\text{idea}}=.39$ of trials ($[.36,.43]$), so source memory accompanied about two out of five detections. When the source was forgotten, participants guessed “AI” with probability $\beta_{\text{idea}}=.43$ ($[.39,.47]$), indicating a bias toward guessing noAI (the HDI excludes .50.50). The false-alarm rate for lure ideas was $f_{\text{idea}}=.29$ ($[.26,.33]$), reflecting a somewhat liberal old–new criterion.
4.4.2. Elaboration component
Elaborations yielded the same high hit rate as ideas, $d_{\text{elab}}=.86$ ($[.84,.88]$), but with a markedly higher source-memory probability, $s_{\text{elab}}=.62$ ($[.57,.65]$), showing that source was correctly identified on well over half of detections. When source memory failed, participants showed a bias toward the AI: $\beta_{\text{elab}}=.60$ ($[.56,.64]$), with the HDI clearly above .50.50. The false-alarm rate was again $f_{\text{elab}}=.29$ ($[.26,.33]$), closely matching the idea component.
4.4.3. Comparison
Posterior contrasts (Elaboration - Idea) confirmed a sizeable source-memory advantage for elaborations and a stronger AI-guessing bias when memory failed: $\Delta s=.23$ (95% HDI $[.17,.28]$, $\Pr(\Delta s>0)=1$) and $\Delta\beta=.18$ ($[.12,.23]$, $\Pr(\Delta\beta>0)=1$). By contrast, sensitivity and criterion were indistinguishable across components: $\Delta d\approx 0$ and $\Delta f\approx 0$.
4.4.4. Summary
Taken together, the results reveal a clear memory advantage for elaboration sources ($s_{\text{elab}}>s_{\text{idea}}$) and a selective AI guessing bias on elaborations ($\beta_{\text{elab}}>.50$ and $\beta_{\text{idea}}$), while overall recognition sensitivity and response criterion are effectively identical for ideas and elaborations. Accordingly, we interpret AI-related differences as shifts in source monitoring and elevated $\beta_{\text{elab}}$, not as changes in item discriminability.
4.5. Qualitative Results
To contextualise our findings, we examined participants’ feedback, and evaluate potential patterns.
4.5.1. Anticipatory Strategies
[…]
4.5.2. Source Attribution Strategies
At the end of the study, we asked participants “Did you employ any strategies for remembering and deciding which ideas or elaborations you came up with on your own or with AI support?”.
[…]
Some participants also revealed in which cases it was particularly easy or difficult to determine the source. Six participants explained that determining the source was easier for topics which they were already familiar with, e.g. one participant answered “The ones related to ocean plastic I remember mostly being related to me because I am very knowledgeable about that topic, so that made those decisions easier”. Four participants mentioned that recalling their own ideas was easier. One mentioned that the interesting ideas were easier to remember. One participant commented that attributing the source for ideas was more difficult than for elaborations, while another commented that it was easier.
4.6. Summary
We summarise the findings as follows:
  • AI involvement reduced item memory accuracy, dropping from 97.6% without any AI to 87.9% for withAI/withAI. This corresponds to substantially lower odds of correct recognition with AI use – especially at the elaboration stage (94.7%, OR = 0.30), compared to a smaller reduction for AI-ideation (97.3%, OR = 0.61).
  • Source memory is supported by consistent workflows: For remembering idea sources, accuracy was highest without any AI (92.4%), followed by withAI/withAI (79.3%), whereas it collapsed in the mixed workflows (withAI/noAI: 37.7%; OR = 0.05, noAI/withAI: 64.0%; OR = 0.14). Remembering elaboration sources showed a similar yet less extreme pattern: noAI/noAI (91.5%) and withAI/withAI (89.2%) outperformed mixed cases (withAI/noAI 79.0%, OR = 0.35; noAI/withAI 80.0%, OR = 0.37). We validated these patterns with the MPT model.
  • AI involvement reduced confidence: Idea source confidence dropped from 85.2 without any AI down to 78.9 for withAI/withAI. Costs were ca. 5-6 points when AI entered at either stage. Elaboration source confidence also dropped, from 86.3 without any AI to around 83 across AI-involved workflows. Metacognitive sensitivity (being more confident when correct) was weak for idea attributions in mixed workflows, but strong for elaboration attributions, with the highest in withAI/withAI.
[…]
5. Discussion
In this study, we explore how well people remember the sources of ideas and texts when working with AI, addressing our research questions:
Our findings show negative effects of AI (RQ1): Recognition of having worked on a solution a week prior was generally high, but AI involvement reduced it. Source attribution for both ideas and elaboration texts was also highest without any use of AI. In summary, using AI reduced people’s ability to remember sources and outcomes of the creative work. We term this effect the “AI Memory Gap.”
Moreover, source attribution accuracy was lowest in mixed workflows (AI support for either ideation or elaboration), indicating that keeping workflows with AI consistent supports memory – although this still did not reach accuracies of not using AI (RQ2).
MPT modelling confirmed that the effects were not due to diminished sensitivity to old items. It confirmed weaker source memory for ideas than elaborations, and revealed different guessing biases when sources were forgotten—towards self for ideas and towards AI for elaborations.
Finally, confidence ratings mirrored these patterns: They were highest without any AI. People overall gave rather high absolute confidence ratings, even when being wrong, and slightly overestimated their overall performance at the end. This indicates that they tended to be overconfident about their own performance (RQ1, RQ3).
5.1. Implications
Following van Berkel and Hornbæk (2023), we identify four key implications of our work for HCI research: theory, design, community, and practice.
5.1.1. Theory
Studies using the Source Monitoring Framework (SMF) (Mitchell and Johnson, 2009) are typically conducted with fixed experimental stimuli. To our knowledge, this work is the first to practically apply the SMF to interaction with generative AI. Our results show that AI use for ideation and writing affects source monitoring, highlighting both the possibility and the importance of updating SMF theory and methodology to such human-AI interaction settings. Concretely, the SMF posits that people attribute mental content to a source by evaluating features of the memory trace (e.g., perceptual details, spatio-temporal context, affect, and records of cognitive operations). Our findings motivate further investigation and extending these in the context of interaction with AI.
5.1.2. Design
These theoretical insights have clear implications for design. AI outputs often lack the perceptual and contextual cues that the SMF identifies as essential for accurate source attribution, making it especially easy to misremember their origin. To directly support memory in human-AI co-creative processes, interaction and UX design should embed salient cues that help users retain a “memory of self-generation.” For example, designs might explore using varying animations, colours, interaction modes, etc. during content generation, which are distinct between own creation vs AI use.
Alternatively, users could be supported in tracking interactions, as demonstrated by the prompt history by Hoque et al. (2024). This could empower users to later verify or reconstruct their own creative process to preserve the memory of who contributed what. In contexts where accurate source attribution is important, it could become critical to design such systems, where both the static text and process traces are part of the final output.
5.1.3. Community
Our work also helps the HCI community contextualize prior research on human-AI interaction, particularly around factors such as responsibility and accountability.
For example, it extends the findings of He et al. (2025b) by showing that disputes over credit in human-AI co-creation may not only stem from social judgments of fairness, but also from systematic memory failures that make accurate source attribution difficult, especially in mixed workflows.
The uncovered “AI memory gap” also provides a psychological lens for connecting with explicit source attribution effects such as the AI Ghostwriter Effect identified by Draxler et al. (2024): Even when users privately feel little ownership of AI generated content, they nonetheless refrain from publicly declaring AI as the author of their creations. Our findings work in parallel as people who genuinely misremember also fail to distinguish the origin of AI-generated ideas. Taken together, both effects can contribute to incorrect attributions that blur the line between accidental and intentional failures to declare AI authorship.
Furthermore, our cognitive perspective adds empirical support for transparency frameworks and themes like those concerning AI disclosure obligations raised by El Ali et al. (2024). If people naturally misremember AI involvement, then voluntary disclosure is unreliable, especially after a longer time gap. Our findings therefore suggest that disclosure obligations cannot depend only on user intent or recollection, but may require transparent, system-supported interaction logging. Ultimately, recognizing and mitigating these memory failures is essential for designing responsible co-creation processes in human-AI interaction.
5.1.4. Practice
Practitioners working in environments where correct source attribution is critical (e.g., creative writing, education, medicine, and law) should be made aware that relying solely on memory can be misleading. This is especially important in the context of our findings on perceived performance: People working with AI tend to overestimate their ability to correctly attribute the source of their creation. As such, end-users should be supported to critically evaluate own recollections in their collaborative human-AI work.
5.2. Principles of Human-AI Interaction in the Context of the AI Memory Gap
Grounded in our results and their implications, we distil a set of principles to guide the design, evaluation, and practical use of AI in human-AI interaction.
Principle of Explicit Source Attribution
As AI involvement systematically impairs source memory, people should not solely (have to) rely on recollection when attributing the source of content. As such, designers should make authorship transitions visible, evaluators should assume that memory-based self-reports are unreliable, and practitioners should adopt strategies that externalise authorship rather than depending on their own memory. Concretely, designers could implement revision histories that highlight which passages were written by a human or suggested by an AI system. Evaluators such as teachers or reviewers could request that submissions include these logs alongside the final product. Practitioners might adopt habits of annotating drafts or saving intermediate snapshots to mark when and how AI was used, in a level of detail that goes beyond merely marking text that was fully AI generated, but also to record where underlying ideas and concepts originated from.
Principle of Consistency
Our findings show that consistent workflows support more accurate memory. Systems where correct attribution might be as important as the final output should therefore support consistent interactions. Concretely, users should be free to involve AI at every stage of the creative process, but they should not be forced or implicitly nudged to rely on AI for any particular task or subtask. When source attribution matters, users may choose to complete each sequence from ideation through elaboration either entirely on their own or entirely with AI. Consistency could also be established by always delegating a specific stage to AI (e.g., always ideate together with AI), while completing the remaining stages alone. And finally, evaluators should recognise that irregular processes are more prone to memory errors.
Principle of Calibrated Confidence
Evaluators should be cautious about interpreting confident statements after using AI systems as evidence of accuracy, as it does not always track accuracy, in particular for ideas in mixed workflows. On the flip side, users should adopt habits of double-checking attributions after having used AI in their workflows and not only trust on their own feelings.
Principle of Cognitive Visibility
While some systems make AI contributions more transparent (e.g., Hoque et al. (2024)), contributions to human thought processes are far harder to capture and later recall. Designers might experiment with tools that record prompts and drafts together, creating a map of thoughts alongside the evolving ideas and making AI’s influence on cognitive processes more visible. However, evaluators should acknowledge that source attributions will always be incomplete and avoid demanding certainty where it cannot exist. Practitioners should be aware that AI may shape not only their writing but also their thinking, often in ways that are difficult to recognize, even for the author themselves.
5.3. Effective Memory Engagement
Here, we dissect our results in light of potential concerns that correct answers could arise without genuine memory engagement. Asking about presented information inadvertently comes with possibilities for answering (correctly) even without actually remembering – in our case by (1) guessing or (2) inferring the answer from cues in the presented text (e.g. “AI-typical” phrasing). We discuss both these “alternative pathways” below.
5.3.1. Memory vs Guessing
Participants performed well above the guessing baseline of 50% in all cases, except for idea source attribution in the withAI/noAI workflow, where they were systematically worse. Moreover, confidence tracked accuracy in most cases, whereas guessing typically produces low subjective certainty (Siedlecka et al., 2021). Thus, our results clearly show that people did not attempt to solve the tasks mainly by guessing.
5.3.2. Memory vs Text Cues
Participants only attributed sources after identifying an item as seen. Already for this first step, any use of AI had a significantly negative influence, indicating that memory formation was affected in Phase 1.
[…]
To dissect this further, the MPT model is specifically designed to isolate item memory, source memory, and guessing. It showed that, after correctly detecting an item as seen, participants recalled the true source without guessing in 39% of cases for ideas and 62% for elaborations. While text cues – which are likely more directly indicative for elaborations – might contribute to the difference between these two values, this does not take away from our result of the “AI memory gap”: That is, isolated from guessing in this way, and combined with the results on accuracy, we see that even with potential text cues, source memory for both idea and elaboration was far from perfect – and significantly worse with AI.
5.4. Limitations, Generalizability and Future work
This study comes with several limitations:
First, we only examined the impact of AI usage during ideation and elaboration tasks as part of the writing process. We opted for those tasks as they represent fundamental stages that precede and influence many other creative processes. Whether our findings apply to other domains, such as coding, image and video creation, or design, needs to be explored in future work. Related, our experimental task mimicked a co-creative workflow for all AI conditions, where the human always made the final decision. Situations in which AI makes executive decisions on its own remain open for investigation. Based on our results, we hypothesise that this would further reduce people’s ability to remember parts of the creative process.
Furthermore, our study focused on a small subset of possible AI interactions, specifically using an LLM (GPT-4o) through a chatbot interface. Using a different type of AI, a different model, or an alternative mode of interaction, such as inline suggestions, could lead to different outcomes.
In addition, people in our study created ideas in one session, whereas real-world creative workflows often stretch over multiple sessions. We also tested people’s source memory after a fixed time gap of approximately one week, and it remains unclear how shorter or longer retention intervals might influence memory performance in this context.
6. Conclusion
Our study shows an “AI memory gap”: Working with AI systematically impairs people’s ability to accurately recall the source of their creative contributions. Memory for authorship was strongest in fully human workflows and weakest in mixed human–AI settings, where mismatched sources disrupted attribution the most. While participants often felt confident in their performance, this confidence exceeded their actual accuracy. Our findings underline the risks of misattribution in human–AI collaboration and call for system designs that make authorship more transparent, as well as further research into how different AI roles, modalities, and interaction settings shape memory and responsibility in creative work.
Acknowledgements.
This project is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 525037874. This work relates to the upcoming ERC project AmplifAI (grant agreement No. 101217557). Daniela Fernandes is funded by the Finnish Doctoral Program Network in Artificial Intelligence, AI-DOC (decision number VN/3137/2024-OKM-6).
Appendix A Appendix
A.1. Problem Statements
  1. (1)
    Come up with simple actions or programs a company can take to help workers spot and avoid scam emails.
  2. (2)
    Think of features that could help blind or low-vision people use a city’s bus/train app more easily.
  3. (3)
    What rules or actions could social media platforms or governments take to slow down or stop fake videos from being shared?
  4. (4)
    What practical things can a family doctor do to help older adults remember and stick to their medicine routine?
  5. (5)
    Come up with ways students can fairly distribute work for an essay in a group project.
  6. (6)
    Oceans are heavily polluted with plastic waste. What practical approaches could help reduce plastic pollution in the oceans?
  7. (7)
    Many countries face demographic challenges with a growing elderly population and fewer young workers. What approaches could help ensure the sustainability of pension funds?
  8. (8)
    Climate change is causing more frequent and severe weather events. What innovative ideas could help communities better prepare for and adapt to these changes?
  9. (9)
    With the rise of remote work, what measures can companies take to maintain employee engagement and productivity outside a traditional office?
  10. (10)
    Come up with practical solutions on how to stay cool during hot summer days.
The known-topic and unknown-topic distractors can be found in our project repository (https://osf.io/anonymous/).
A.2. Questionnaire Questions
  1. (1)
    How many of the ideas (in percent) do you think you assigned correctly? (0-100)
  2. (2)
    Do you think you assigned more ideas incorrectly to AI or to yourself?
    • More incorrectly assigned to AI
    • More incorrectly assigned to myself
    • About equally to both
  3. (3)
    How many of the solution texts (in percent) do you think you assigned correctly? (0-100)
  4. (4)
    Do you think you assigned more solution texts incorrectly to AI or to yourself?
    • More incorrectly assigned to AI
    • More incorrectly assigned to myself
    • About equally to both
  5. (5)
    Did you employ any strategies for remembering and deciding which ideas or elaborations you came up with on your own or with AI support?
  6. (6)
    Gender:
    • Male
    • Female
    • Non-binary
    • Prefer not to say
    • Prefer to self describe:
  7. (7)
    English Proficiency:
    • Native speaker
    • Fluent (C2)
    • Advanced (C1)
    • Upper-intermediate (B2)
    • Intermediate (B1)
    • Pre-intermediate (A2)
    • Beginner (A1)
  8. (8)
    Highest Degree:
    • PhD/Doctorate
    • Master’s degree
    • Bachelor’s degree
    • Associate degree
    • High school diploma
    • Some college (no degree)
    • Other
  9. (9)
    How often do you use AI tools (e.g., ChatGPT, Claude, Copilot)?
    • Daily
    • Several times a week
    • About once a week
    • A few times a month
    • Rarely
    • Never
A.3. Misc
Section Deviation
Participant recruitment We recruited participants until reaching a sample of 200 valid and complete responses in Phase 1. To replace the drop-outs, we started 5 sequential recruitment rounds. Due to time-constraints we decided against starting another round because each round would delay data collection by another week. This resulted in the final sample of 184 people who successfully completed both Phases of the study.
Condition labels We exchanged the condition label as used in this paper, using “withAI” instead of “AI”, to clarify that ideas and elaborations were co-created by the participants together with AI, rather than generated automatically by AI alone. The label for the condition without any AI involvement remained “noAI”.
Table 9. Rationale for deviations from the pre-registration. The pre-registration can be accessed at: https://aspredicted.org/g78p-w7db.pdf
[…]
The table contains two rows rationale for deviations from the pre-registration.